OpenStudy (anonymous):

What slope passes through (5, 5) and (-4, 5)???

OpenStudy (zzr0ck3r):

(y_2-y_1)/(x_2-x_1) = (5-5)/(-4-5) = ?

OpenStudy (anonymous):

0/-9?

OpenStudy (zzr0ck3r):

or think about what a slope would have to be, if the function is linear, for it to pass through the same y point (5) more than once.

OpenStudy (zzr0ck3r):

0/-9 = ?

OpenStudy (anonymous):

9

OpenStudy (zzr0ck3r):

what is 0/1?

OpenStudy (anonymous):

-1?

OpenStudy (zzr0ck3r):

0/anything = 0

OpenStudy (zzr0ck3r):

unless its 0/0 but dont worry about that

OpenStudy (zzr0ck3r):

so 0/1 = 0 0/2 = 0 0/9 = 0 whats 0/-9 = ?

OpenStudy (zzr0ck3r):

@jdillon ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

now can you help me find the slope of this -4,1 and -2,-3

OpenStudy (zzr0ck3r):

remember what slop is rise/run and since the two points have a y value of 5 there is 0 rise and 0 over anything is 0

OpenStudy (anonymous):

so the answer is 0?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

no it shows that its not it

OpenStudy (anonymous):

-4,1 and -2,-3???

OpenStudy (zzr0ck3r):

? you sure its the right question?

OpenStudy (anonymous):

promise

OpenStudy (anonymous):

-4,1 is the rise and -2, -3 is the run

OpenStudy (anonymous):

what about (9, -6) and (5, 10)

OpenStudy (zzr0ck3r):

10--6/5-9 16/-4 = -4

OpenStudy (zzr0ck3r):

the book is wrong then

OpenStudy (zzr0ck3r):

first off the slope will be one number not 4